OpenStudy (anonymous):

What is the height of cylinder with a surface area of 226.08 square meters and a radius of 3 meters? (Use 3.14 for π.) 27 meters 18 meters 3 meters 9 meters

OpenStudy (anonymous):

Well, you know the surface area of the top(and bottom) to be: \[\pi*r^2\] And, the surface area of the outside part of the cylinder is \[2*\pi*r*h\] So, now, our total surface area is \[2*\pi*r*h+\pi*r^2+\pi*r^2\] So, after combining like terms, and such, we should see \[18\pi+6\pi*h=226.08\] Now, solving for h should be some simple algebra. \[h=(226.08-18\pi)/6\pi\] \[h=8.999m\] Or, 9meters.